# Deep Neural networks for solving high-dimensional parabolic partial differential equations

### v1

@@ 2.1 Strategy A: PDE residual based PINN methods
We consider the following elliptic PDE version of (1) (without the time derivative) defined on a domain $\Omega\subset\mathbb{R}^{d}$ with the boundary condition (5) defined on $\Gamma\subset\mathbb{R}^{d}$, i.e.,
@@ 2.2.1 Backward schemes
Solution to the process $Y_{t}$ provided with terminal values can be progressively solved in a dynamical programming (DP) manner, i.e. in backward order of time steps, so that the problem is also progressively resolved. When stepwise DNNs are trained, transfer learning can be applied from the outcome of previous steps to newly initialized networks. The Feynman–Kac formula [oksendal2003stochastic] represents $Y_{t}$ by an expectation conditioned by the natural filtration at time $t$. For quasi-linear PDEs, the evaluation often involves future values of $Y_{s}$ for $s\in(t,T]$, which may be handled stepwisely to reduce approximation error. Huré et al. [Pham2020DBDP] proposed backward DP-type schemes training DNNs that approximate $u(t,\cdot)$ and $\nabla u(t,\cdot)$ on descending time steps, with the loss functions defined as square of the residual of discretized SDE on each time step. Pham et al. [pham2021neural] extended their work to fully nonlinear PDEs. Beck et al. [Beck2021DeepSplitting] developed backward schemes for $u(t,\cdot)$ that are ruled by the $L^{2}$-minimality property of stepwise conditional expectations [klenke2008probability]. As the Feynman-Kac formula can be viewed as a generalized characteristic method connecting the terminal condition to the solution at earlier times (backward). Han et al. [han2024picard] incorporates Monte Carlo simulations of the Feynman–Kac formula [oksendal2003stochastic] for $Y_{t}$ and the Malliavin method [ma2002representation] for $Z_{t}$ for a Picard fix-point iteration scheme for the PDE solution represented by a DNN.
@@ 2.3 Strategy C: Random differences by expectation
When simulation of $Y_{t}$ is implemented with DNNs, the DNN calls accumulate under (usually nonlinear) operations in the scheme, slowing down the backward propagation, thus limiting the scalability for solving very high-dimensional problems or on a large time interval. Such issue may be mitigated by slicing the trajectory into short pieces if localized properties of trajectories can be effectively used. Xu and Zhang [Xu2025Shotgun] showed that a single-step SDE discretization by FBSNN, giving a random difference operator, is an estimator of the residual of the PDE, and proposed a PINN-based scheme that applies coarse time sampling for accessing the approximated distribution of $X_{t}$. Cai et al. [Cai2025RDO] proposed a Galerkin weak form implementation of the random difference operator approach, mathematically equivalent to DeepMartNet.
@@ 4.1 Itô formula and Pardoux-Peng nonlinear Feynman-Kac formula
where the initial state $X_{0}$ is given, and the functions $\mu:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ are the drift and diffusion coefficients, respectively. Let $h:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ be a function of class $C^{1,2}$, i.e., continuously differentiable in $t$ and twice continuously differentiable in $\mathbf{x}$. Then Itô’s formula [weinanbook, oksendal2003stochastic] says that the process $h(t,X_{t})$ is also an Itô process, governed by the forward SDE
@@ 4.4 Random difference and shotgun neural networks
Equation (30) can be viewed as a random difference method (RDM) for the PDE operator as seen from (32). A general approach of deriving such RDMs can be done using simple Taylor expansions [Xu2025Shotgun, Cai2025RDO], which were in fact also used to derive the Itô formula (15) [weinanbook, oksendal2003stochastic].
@@
(41) $D_{h}(t,\textbf{x},\xi;u)=\frac{u(t+h,\textbf{x}+\mu h+\sqrt{h}\sigma\xi)-u(t,\textbf{x})}{h},$
@@
(45) $\bar{\mathcal{D}}_{h}u(t,\textbf{ x}):=\mathbb{E}\left[\frac{D_{h}(t,\textbf{x},\xi;u)+D_{h}(t,\textbf{x},-\xi;u)}{2}\right]$
@@
Finally, we note that the approximation of differential operators by expectation of random differences bears similarity to the approximation (60) (61) given in the Appendix A.3.5 for the RS-PINN with the former having an upwinding nature corresponding to the drift of the system [Cai2025RDO].
@@ 5.2 Black-Scholes (BS) equation
We examine the solution at $t=0$ within $x\in[90,110]^{d}$ and at $t=0$ for $d=100,1000$ at 1000 test points uniformly from the cube [beck2021solving]. The results are summarized in Table 2 with the network structures of each method given in Appendix A.4.2.
For SDE based methods, a scaled version of (49) will be used with the following scaling $\textbf{z}:=\textbf{x}/K$, $v(t,\textbf{z}):=u(K\textbf{ z},t)/K$, $K=100$ and the scaled BS equation takes the exact same form of (49) with x replaced by z and $u$ by $v$ and the exact solution is given by $v(t,\textbf{z})=\mathbb{E}\left[\phi_{1}\left(y_{1},y_{2},\cdots,y_{d}\right)\right].$

### v2

@@ Abstract
keywords:
High dimensional PDE; Neural networks; Curse of dimensionality
{MSCcodes}
49L12, 49L20, 49M05, 65C30, 93E20
@@ 2.1 Strategy A: PDE residual based PINN methods
We consider the following elliptic PDE version of (1) (without the time derivative) defined on a domain $\Omega\subset\mathbb{R}^{d}$ with the boundary condition (5) defined on $\Gamma\subset\mathbb{R}^{d}$,
@@ 2.2.1 Backward schemes
Solution to the process $Y_{t}$ provided with terminal values can be progressively solved in a dynamical programming (DP) manner, i.e. in backward order of time steps, so that the problem is also progressively resolved. When stepwise DNNs are trained, transfer learning can be applied from the outcome of previous steps to newly initialized networks. The Feynman–Kac formula [weinanbook] represents $Y_{t}$ by an expectation conditioned by the natural filtration at time $t$. For quasi-linear PDEs, the evaluation often involves future values of $Y_{s}$ for $s\in(t,T]$, which may be handled stepwisely to reduce approximation error. Huré et al. [Pham2020DBDP] proposed backward DP-type schemes training DNNs that approximate $u(t,\cdot)$ and $\nabla u(t,\cdot)$ on descending time steps, with the loss functions defined as square of the residual of discretized SDE on each time step. Pham et al. [pham2021neural] extended their work to fully nonlinear PDEs. Beck et al. [Beck2021DeepSplitting] developed backward schemes for $u(t,\cdot)$ that are ruled by the $L^{2}$-minimality property of stepwise conditional expectations [klenke2008probability]. As the Feynman-Kac formula can be viewed as a generalized characteristic method connecting the terminal condition to the solution at earlier times (backward). Han et al. [han2024picard] incorporates Monte Carlo simulations of the Feynman–Kac formula [weinanbook] for $Y_{t}$ and the Malliavin method [ma2002representation] for $Z_{t}$ for a Picard fix-point iteration scheme for the PDE solution represented by a DNN.
@@ 2.3 Strategy C: Random differences by expectation
When simulation of $Y_{t}$ is implemented with DNNs, the DNN calls accumulate under (usually nonlinear) operations in the scheme, slowing down the backward propagation, thus limiting the scalability for solving very high-dimensional problems or on a large time interval. Such issue may be mitigated by slicing the trajectory into short pieces if localized properties of trajectories can be effectively used. Xu and Zhang [Xu2025Shotgun] showed that a single-step SDE discretization by FBSNN, giving a random difference operator, is an estimator of the residual of the PDE, and proposed a PINN-based scheme that applies coarse time sampling for accessing the approximated distribution of $X_{t}$. Cai et al. [Cai2025RDO] proposed a Galerkin weak form of the random difference approach, an equivalent to DeepMartNet.
@@ 4.1 Itô formula and Pardoux-Peng nonlinear Feynman-Kac formula
where the initial state $X_{0}$ is given, and the functions $\mu:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ are the drift and diffusion coefficients, respectively. Let $h:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ be a function of class $C^{1,2}$, i.e., continuously differentiable in $t$ and twice continuously differentiable in $\mathbf{x}$. Then Itô’s formula [weinanbook] says that the process $h(t,X_{t})$ is also an Itô process, governed by the forward SDE
@@ 4.4 Random difference and shotgun neural networks
Equation (30) can be viewed as a random difference method (RDM) for the PDE operator as seen from (32). A general approach of deriving such RDMs can be done using simple Taylor expansions [Xu2025Shotgun, Cai2025RDO], which were in fact also used to derive the Itô formula (15) [weinanbook].
@@
(41) $D_{h}(t,\textbf{x},\xi;u)=\left(u(t+h,\textbf{x}+\mu h+\sqrt{h}\sigma\xi)-u(t,\textbf{x})\right)/h,$
@@
(45) $\bar{\mathcal{D}}_{h}u(t,\textbf{ x}):=\frac{1}{2}\mathbb{E}\left[D_{h}(t,\textbf{x},\xi;u)+D_{h}(t,\textbf{x},-\xi;u)\right]$
@@ 5.2 Black-Scholes (BS) equation
We examine the solution at $t=0$ within $x\in[90,110]^{d}$ and at $t=0$ for $d=100,1000$ at 1000 test points uniformly from the cube [beck2021solving]. The results are summarized in Table 2 with the network structures given in Appendix A.4.2. For SDE based methods, a scaled version of (49) is used with the following scaling $\textbf{z}:=\textbf{x}/K$, $v(t,\textbf{z}):=u(K\textbf{ z},t)/K$, $K=100$.